2 cos a cot^2 a - 6 cos a - cot^2 a = -3 is it an identity or a conditional equation? if not an identity, solve for the unknown


1
Expert's answer
2021-05-07T11:45:07-0400

The given equation "2cosa.cot^2a-6cosa-cot^2a=-3" is identity equation.

Solving the above equation,

"2cosa.cot^2a-6cosa-cot^2a=-3\newline
2cot^2a.cosa-cot^2a+3-6cosa=0\newline
cot^2a(2cosa-1)-3(2cosa-1)=0\newline
(cot^2a-3)(2cosa-1)=0\newline
(cota+ \sqrt3)(cota-\sqrt3)(2cosa-1)=0"

Then, we get

"cota+ \sqrt3=0..............................1\newline
cota-\sqrt3=0..............................2\newline
2cosa-1=0................................3"

General solution of 1 is "a=n\pi +\frac{5 \pi }{6}" , 2 is "a=n\pi +\frac{\pi }{6}" and 3 is "a=2n\pi +\frac{\pi }{3} \text{and}\space a=2n\pi +\frac{5\pi }{3}".

Thus, the required solution are "a=n\pi +\frac{5 \pi }{6}", "a=n\pi +\frac{\pi }{6}", "a=2n\pi +\frac{\pi }{3} \text{and}\space a=2n\pi +\frac{5\pi }{3}".